# Dark Dwarfs: Dark Matter-Powered Sub-Stellar Objects Awaiting Discovery at the Galactic Center

### v1

@@ Abstract
Discovering the microphysical description of dark matter (DM) is a paramount goal of particle physics and cosmology. All of our current evidence for DM’s existence is indirect, and we have no information pertaining to its mass and interactions with visible matter or other dark sector particles. We do however know that DM needs to be produced in the early universe. Among the many mechanisms for this, thermal freeze-out is a well-motivated minimal scenario. Provided that Standard Model (SM) particles are present in the final states of the process that sets the relic density, there is an energy transfer to the visible sector Steigman and Turner (1985); Jungman et al. (1996); Bertone et al. (2005); Bertone and Hooper (2018). This process can be detected via annihilation signals in space Geringer-Sameth et al. (2015); Ackermann et al. (2015), or scattering in the laboratory Smirnov and Beacom (2020); Parikh et al. (2023), but also affects astrophysical objects Batell et al. (2010); Pospelov et al. (2008); Rothstein et al. (2009); Pospelov and Ritz (2009); Chen et al. (2009); Schuster et al. (2010a, b); Bell and Petraki (2011); Feng et al. (2016a); Kouvaris and Tinyakov (2010a); Feng et al. (2016b); Allahverdi et al. (2017); Leane et al. (2017); Arina et al. (2017); Albert et al. (2018a, b); Nisa et al. (2019); Niblaeus et al. (2019); Cuoco et al. (2020); Serini et al. (2020); Acevedo et al. (2021); Mazziotta et al. (2020); Bell et al. (2021); Bose et al. (2022a); Smirnov et al. (2024); Croon and Sakstein (2024); John et al. (2024).
Objects that are not supported by nuclear burning such as white dwarfs, neutron stars Goldman and Nussinov (1989); Gould et al. (1990); Kouvaris (2008); Bertone and Fairbairn (2008); de Lavallaz and Fairbairn (2010); Kouvaris and Tinyakov (2010b); McDermott et al. (2012); Kouvaris and Tinyakov (2011a); Guver et al. (2014); Bramante et al. (2013); Bell et al. (2013); Bramante et al. (2014); Bertoni et al. (2013); Kouvaris and Tinyakov (2011b); McCullough and Fairbairn (2010); Angeles Perez-Garcia and Silk (2015); Bramante (2015); Graham et al. (2015); Cermeno et al. (2016); Graham et al. (2018); Acevedo and Bramante (2019); Janish et al. (2019); Krall and Reece (2018); McKeen et al. (2018); Baryakhtar et al. (2017); Raj et al. (2018); Bell et al. (2018); Chen and Lin (2018); Garani et al. (2019); Dasgupta et al. (2019); Hamaguchi et al. (2019); Camargo et al. (2019); Bell et al. (2019); Acevedo et al. (2020); Joglekar et al. (2019, 2020); Bell et al. (2020); Garani et al. (2020); Leane et al. (2021); Bose et al. (2022b); Collier et al. (2022), planets Mack et al. (2007); Bramante et al. (2020); Leane and Linden (2021); Croon and Smirnov (2023), and brown dwarfs Leane and Smirnov (2021); Leane et al. (2021) have been shown to exhibit observable signatures based on DM interactions, as an additional energy injection can lead to changes in their structure and evolution. These objects are also excellent probes of gravity Sakstein (2015a, b).
@@
Figure 1: Effects of DM energy injection on luminosity and lithium depletion for objects with masses $M_{0}\equiv M/M_{\odot}$pper panel: time-evolution of the total luminosity $L_{\rm surf}$ (solid lines), luminosity from DM burning $L_{\rm DM}$ (dashed lines) and luminosity in hydrogen burning (dotted lines). Lower panel: time evolution of the surviving fraction of 7Li. Dark dwarfs retain their initial lithium-7 while red dwarfs deplete it.
@@
In this work we investigate the effects of energy injection from DM annihilation on celestial bodies at the stellar mass boundary. These objects, which have masses $M\lesssim 0.1{\rm M}_{\odot}$, are too light to fuse hydrogen on the PPI chain due to their cool core temperatures $T\sim 10^{6}$ K. Those heavier than $\sim$0.$0.075{\rm M}_{\odot}$, red dwarfs or M-dwarfs, evolve into equilibrium states that fuse hydrogen to helium-3.  Lighter bodies, so called brown dwarfs, are unable to do so. They experience transitory periods of hydrogen, lithium, and deuterium burning, but ultimately cool and contract eternally.
@@
As shown in Fig. 1 the situation is dramatically different when DM annihilation is present. Contracting objects attain an equilibrium configuration where they are supported primarily by DM heating. This steady-state exists for any non-zero DM density, but is attained in a Hubble time for $\rho_{\rm DM}\gtrsim 10^{3}\rm GeV\,cm^{-3}$. We name these objects Dark Dwarfs (DDs), and differentiate them from red dwarfs near the hydrogen burning limit — $M\sim 0.075{\rm M}_{\odot}$ in the SM — which are supported by a combination of nuclear burning and DM heating. DDs are physically distinct from brown/red dwarfs in several ways:
@@
These properties, which are exemplified in Fig. 1, are independent of the DM mass but do depend upon the ambient DM density and velocity dispersion. DDs can be identified by their enhanced lithium abundance despite a relatively large mass, and old stellar age.
Analytic Model for Dark Dwarfs
Sub-stellar objects are relatively simple compared with more massive bodies. Their homogeneous structure and lack of strong nuclear burning make them well-described by analytic models. They are supported by a combination of the degeneracy pressure of the electrons and the gas pressure of the ions. Under these assumptions the equation of state (EOS) is that of an $n=3/2$ polytrope except near the surface where there is a transition to a phase of molecular hydrogen Auddy et al. (2016). The full details of this model are reviewed in the Supplementary Material.
@@
$\frac{L_{\rm surf}}{L_{\odot}}=\frac{0.0578353b_{1}^{3}M_{0}}{\kappa_{R}}\psi^{3\nu}\left(\frac{M_{0}^{5/3}\psi^{-\nu}}{b_{1}\kappa_{R}(1+\gamma+\alpha\psi)^{2}}\right)^{1/7}.$ (1)
where $M_{0}=M/M_{\odot}$, $\kappa_{R}$ is the Rosseland mean opacity at the photosphere (we set $\kappa_{R}=0.01$ cm2/g in what follows, typical for these objects), $b_{1}$ and $\nu$ are parameters of the molecular hydrogen EOS (we shall use parameter point D in the numerical calculations in this work, defined in table S1 in the Supplemental Material, with $b_{1}=2$ and $\nu=1.6$), $\psi=k_{B}T/E_{F}$ is the degeneracy parameter, and $\gamma$ is a function of $\psi$ given in Eq. (S4). The properties of the star are completely determined once a value of $\psi$ is specified. The steady-state condition that determines $\psi$ is energy conservation — the surface luminosity must balance the luminosity from all heat sources:
$L_{\rm surf}=\sum_{i}L_{i}$ (2)
where $i$ runs over all processes injecting energy into the star, which in our case include nuclear burning and dark matter annihilation. The polytropic approximation ceases to be valid in very low mass objects because Coulomb scattering begins to become important Clayton (1983); Saumon et al. (1996); Burrows and Liebert (1993). This breakdown happens for $M\lesssim 0.18{\rm M}_{\odot}$ in the SM and, as derived in Appendix I.4, breaks down at lower masses when the DM energy injection is important. The $n=3/2$ approximation remains valid for all objects studied in this work.
@@
[…]
The luminosity from these processes, assuming the $n=3/2$ polytrope, is Auddy et al. (2016)
$\frac{L_{\rm HB}}{L_{\odot}}=7.33\times 10^{16}M_{0}^{11.977}\frac{\psi^{6.0316}}{(1+\gamma(\psi)+\alpha\psi)^{16.466}},$ (5)
$M_{0}=0.03362\,b_{1}^{0.26605}\frac{(1+\gamma(\psi)+\alpha\psi)^{1.5069}}{\psi^{0.266053\nu-0.56165}}\,.$ (6)
There is a minimum value of $\psi$ for which Eq. (6) has solutions. Below this, the object cannot burn hydrogen stably because the core temperatures and densities are too low to sustain a sufficient nuclear burning rate to balance the surface luminosity. This boundary corresponds to the minimum mass for hydrogen burning (MMHB). Stars heavier than this are red dwarfs while lighter objects are brown dwarfs. In our model, the specific value is $M_{\rm MMHB}=0.075{\rm M}_{\odot}$, consistent with the literature e.g., Burrows and Liebert (1993); Auddy et al. (2016).
[…]
@@
$\begin{split}L_{\rm DM}&=m_{\rm DM}\pi R^{2}\Phi,\quad\text{with}\\ \Phi&=\sqrt{\frac{8}{3\pi}}\frac{\rho_{\rm DM}v_{\rm DM}f_{\rm cap}}{m_{\rm DM}}\left(1+\frac{3}{2}\frac{v_{\rm esc}^{2}}{v_{\rm DM}^{2}}\right)\end{split}$ (7)
@@
where $v_{\rm esc}^{2}=2GM/R$ is the escape velocity, $\rho_{\rm DM}$ is the ambient DM density, $v_{\rm DM}$ is circular velocity of galactic DM, $f_{\rm cap}$ is the fraction of DM captured that we set to unity, and $m_{\rm DM}$ is the DM mass; the term in the brackets accounts for gravitational focusing. Throughout this work, we fix $v_{\rm DM}=50$ km/s as a conservative benchmark Karukes et al. (2020). These equations have an important consequence: the DM luminosity is independent of the DM mass. Since the properties of the object are determined solely by energy conservation, the characteristics of objects supported partially or fully by DM annihilation are similarly DM mass-independent.
@@
To understand the differences between hydrogen and DM burning, it is enlightening to consider a star supported entirely by DM annihilation, i.e., $L_{\rm DM}=L_{\rm surf}$. Using equation (S5) for the radius in (7), we can find an analytic solution when gravitational focusing is neglected:
@@
$\begin{split}M_{0}=&8.58\times 10^{-7}\times\frac{(1+\gamma(\psi)+\alpha\psi)^{1.19895}}{\psi^{1.4993\nu}}\\ &\times\left(\frac{\rho_{\rm DM}}{\rm GeV\,cm^{-3}}\frac{v_{\rm DM}}{50\rm km\,s^{-1}}\frac{f_{\rm cap}}{b_{1}^{2.857}}\right)^{0.52475}.\end{split}$ (8)
@@
In contrast to the equivalent formula for hydrogen-supported objects (6), Eq. (8) has no minimum — DM can support structures of arbitrarily small masses. This is because, unlike hydrogen burning, the DM burning rate does not depend on the stellar properties — DM burning is ever-present. Including gravitational focusing does not alter this conclusion and in fact produces stronger effects. It is included in our numerical studies below.
@@
Time Evolution and Stability
To study the time evolution that leads to the formation of dark dwarfs, we use the cooling model introduced in Stevenson (1991); Burrows and Liebert (1993). Using the first law of thermodynamics, the equation for the energy of a contracting star is
$T\frac{\mathrm{d}S}{\mathrm{d}t}=\varepsilon-\frac{\partial L_{\rm surf}}{\partial M}$ (9)
where $S$ is the entropy per unit mass and $\varepsilon$ is the energy generation rate per unit mass from burning. Integrating this equation, employing the equation of state for molecular hydrogen, and using the EOS and properties of $n=3/2$ polytropes one finds Auddy et al. (2016)
$\frac{\mathrm{d}\psi}{\mathrm{d}t}=\frac{\bar{\mu}}{1.5\mu_{e}^{\frac{8}{3}}E_{0}}M_{0}^{{-7/3}}(1+\gamma(\psi)+\alpha\psi)^{2}(L_{\rm burn}-L_{\rm surf}),$ (10)
@@
$\frac{1}{\bar{\mu}}=\frac{1}{{\mu}}+\frac{3X_{H^{+}}(1-X_{H^{+}})}{2(2-X_{H^{+}})}$ (11)
@@
with $X_{H^{+}}$ the mass fraction of ionized hydrogen; $\bar{\mu}=1.02$ for parameter point D, see the Supplementary Material. In the absence of any burning, equation (10) implies that $\psi$ will decrease from its initial value i.e., that over time $k_{B}T$ will decrease relative to $E_{F}$ (see equation (S3)) and the star will be increasingly supported by degeneracy pressure. In the presence of burning, there is a steady-state when the right hand side of equation (10) is zero i.e., when $L_{\rm burn}=L_{\rm surf}$. These are the solutions we derived above. We thus expect that any initial configuration will evolve towards this steady-state, reaching it at sufficiently late times.
@@
The stability of steady-state solutions can be determined as follows. Letting $\psi_{0}$ be the value of $\psi$ where the steady-state is achieved, we can write $\psi(t)=\psi_{0}+\delta\psi(t)$ and Taylor-expand equation (10) to find $\dot{\delta\psi}=f(\psi_{0})\delta\psi+\mathcal{O}(\delta\psi^{2})$, where $f(\psi_{0})$ is the derivative of the right hand side of (10) with respect to $\psi$ and evaluated at $\psi_{0}$. If $f(\psi_{0})<0$ then the equilibrium is stable. The resulting expressions are long and not informative so we do not give them here. They can be found in our accompanying  code. Using the same code, we found that dark dwarfs are always stable. This is true for objects supported by a combination of H-burning and DM burning, or objects supported solely by DM burning. In contrast, only the branch of SM stars (solely H-burning) with $\psi>\psi_{\rm min}$ is stable.
@@
Figure 3: Time evolution of important stellar properties. The continuous lines correspond to objects evolving in the presence of DM annihilation while dashed lines indicate the SM predictions. Four representative stellar masses indicated in the top panel are shown.
@@
Figure 4: Time evolution of the core temperatures of five objects with different masses with (solid lines) and without (dashed lines) DM energy injection. Coulomb corrections to the pressure hare included.
@@
In Fig. 3 we show properties of objects in the presence of DM annihilation as a function of time found by solving Eq. (10).  Several differences between SM objects and those with some amount of DM burning are evident. SM brown dwarfs lighter than the MMHB cool continuously, contracting and becoming dimmer. SM objects heavier than the MMHB cool and evolve to red dwarfs where they are supported by H-burning. In contrast, in the presence of DM burning lighter objects cool similarly to brown dwarfs until they reach the steady-state dark dwarf solution with constant core temperature, radius, and brightness. The time to reach the dark dwarf state is shorter for larger ambient DM densities. The MMHB is also DM density dependent.
@@
Lithium burning
The lithium test is a primary method for confirming that an object is a brown dwarf Rebolo et al. (1992). Observing lithium lines in stellar spectra allows astronomers to trace the core temperature history of young stars and brown dwarfs, distinguishing between different evolutionary epochs. Therefore, deviations from the SM expectation of the lithium-7 abundance may provide a method for detecting dark dwarfs.

### v2

@@ 1 Introduction
Discovering the microphysical description of dark matter (DM) is a paramount goal of particle physics and cosmology. All of our current evidence for DM’s existence is indirect, and we have no information pertaining to its mass and interactions with visible matter or other dark sector particles. We do however know that DM needs to be produced in the early universe. Among the many mechanisms for this, thermal freeze-out is a well-motivated minimal scenario. Provided that Standard Model (SM) particles are present in the final states of the process that sets the relic density, there is an energy transfer to the visible sector [1, 2, 3, 4]. This process can be detected via annihilation signals in space [5, 6], or scattering in the laboratory [7, 8], but also affects astrophysical objects [9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35].
Objects that are not supported by nuclear burning such as white dwarfs, neutron stars [36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76], planets [77, 78, 79, 80], and brown dwarfs [81, 74] have been shown to exhibit observable signatures based on DM interactions, as an additional energy injection can lead to changes in their structure and evolution. Note that large fractions of the dark matter parameter space can lead to substantial capture rates in celestial bodies. For example, in the case of spin-dependent proton scattering, cross sections above $\sim 10^{-25}\,\rm{cm}^{2}$ are practically unconstrained [82], with some exceptions which depend on modeling assumptions e.g., cosmology dependent constraints from the CMB [83], the lack of MW satellites [84] for velocity-independent scattering cross sections, and the existence of Jupiter [80]. In particular, the mass range in the GeV region at large cross sections is not constrained by Earth and Mars heating [78] because evaporation would eject dark matter particles in that mass range from rocky planets but would retain them in gas giants and brown dwarfs [85]. Such cross-sections would lead to geometric capture rates in brown dwarfs, containing predominantly hydrogen nuclei. Thus, masses above the evaporation threshold at the GeV scale are testable in heating scenarios. Note that these objects are also excellent probes of modified theories of gravity [86, 87].
@@
Figure 1: Effects of DM energy injection on luminosity and lithium depletion for objects with masses $M_{0}\equiv M/M_{\odot}$ embedded in DM density $\rho_{0}\equiv\rho_{\rm DM}/(\rm GeV\,cm^{-3})$Upper panel: time-evolution of the total luminosity $L_{\rm surf}$ (solid lines), luminosity from DM burning $L_{\rm DM}$ (dashed lines) and luminosity in hydrogen burning (dotted lines). Lower panel: time evolution of the surviving fraction of 7Li. Dark dwarfs retain their initial lithium-7 while red dwarfs deplete it.
@@
In this work we investigate the effects of energy injection from DM annihilation on celestial bodies at the stellar mass boundary. These objects, which have masses $M\lesssim 0.1{\rm M}_{\odot}$, are too light to fuse hydrogen on the PPI chain due to their cool core temperatures $T\sim 10^{6}$ K. Those heavier than $\sim$$0.075{\rm M}_{\odot}$, red dwarfs or M-dwarfs, evolve into equilibrium states that fuse hydrogen to helium-3.  Lighter bodies, so called brown dwarfs, are unable to do so. They experience transitory periods of hydrogen, lithium, and deuterium burning, but ultimately cool and contract eternally.
@@
As shown in Fig. 1 the situation is dramatically different when DM annihilation is present. Contracting objects attain an equilibrium configuration where they are supported primarily by DM heating. This steady-state exists for any non-zero DM density, but is attained in a Hubble time for $\rho_{\rm DM}\gtrsim 10^{3}\rm GeV\,cm^{-3}$ (for reference, assuming an NFW profile, the DM density a parsec away from the Galactic Centre (GC) is $\rho_{\rm DM}=10^{4}\,\rm GeV\,cm^{-3}$). We name these objects Dark Dwarfs (DDs), and differentiate them from red dwarfs near the hydrogen burning limit — $M\sim 0.075{\rm M}_{\odot}$ in the SM — which are supported by a combination of nuclear burning and DM heating. DDs are physically distinct from brown/red dwarfs in several ways:
@@
These properties, which are exemplified in Fig. 1, are independent of the DM mass but do depend upon the ambient DM density and velocity dispersion. Moreover, through our novel analytic model we demonstrate that DM heating leads to non-linear feedback that affects the energy emission for objects near the $0.075{\rm M}_{\odot}$ boundary, as can be seen in Fig. 1. A consequence of this is that DDs can be identified by their enhanced lithium abundance despite a relatively large mass, and old stellar age.
2 Analytic Model for Dark Dwarfs
Sub-stellar objects are relatively simple compared with more massive bodies. Their homogeneous structure and lack of strong nuclear burning make them well-described by analytic models. They are supported by a combination of the degeneracy pressure of the electrons and the gas pressure of the ions. Under these assumptions the equation of state (EOS) is that of an $n=3/2$ polytrope except near the surface where there is a transition to a phase of molecular hydrogen [88]. The full details of this model are reviewed in the Supplementary Material.
@@ 2 Analytic Model for Dark Dwarfs
$\frac{L_{\rm surf}}{L_{\odot}}=\frac{0.0578353b_{1}^{3}M_{0}}{\kappa_{R}}\psi^{3\nu}\left(\frac{M_{0}^{5/3}\psi^{-\nu}}{b_{1}\kappa_{R}(1+\gamma+\alpha\psi)^{2}}\right)^{1/7}.$ (2.1)
where $M_{0}=M/M_{\odot}$, $\kappa_{R}$ is the Rosseland mean opacity at the photosphere (we set $\kappa_{R}=0.01$ cm2/g in what follows, typical for these objects), $b_{1}$ and $\nu$ are parameters of the molecular hydrogen EOS (we shall use parameter point D in the numerical calculations in this work, defined in table 1 in the Supplemental Material, with $b_{1}=2$ and $\nu=1.6$), $\psi=k_{B}T/E_{F}$ is the degeneracy parameter, and $\gamma$ is a function of $\psi$ given in Eq. (A.4). The properties of the star are completely determined once a value of $\psi$ is specified. The steady-state condition that determines $\psi$ is energy conservation — the surface luminosity must balance the luminosity from all heat sources:
$L_{\rm surf}=\sum_{i}L_{i}$ (2.2)
where $i$ runs over all processes injecting energy into the star, which in our case include nuclear burning and dark matter annihilation. The polytropic approximation ceases to be valid in very low mass objects because Coulomb scattering begins to become important [89, 90, 91]. This breakdown happens for $M\lesssim{0.018}{\rm M}_{\odot}$ in the SM and, as derived in Appendix A.4, breaks down at lower masses when the DM energy injection is important. The $n=3/2$ approximation remains valid for all objects studied in this work.
@@
$\textrm{p}+\textrm{p}$ $\rightarrow\textrm{d}+\textrm{e}^{+}+\nu_{e}$ (2.3)
$\textrm{p}+\textrm{d}$ $\rightarrow^{3}\!\!{\rm He}+\gamma.$ (2.4)
The luminosity from these processes, assuming the $n=3/2$ polytrope, is [88]
$\frac{L_{\rm HB}}{L_{\odot}}=7.33\times 10^{16}M_{0}^{11.977}\frac{\psi^{6.0316}}{(1+\gamma(\psi)+\alpha\psi)^{16.466}},$ (2.5)
$M_{0}=0.03362\,b_{1}^{0.26605}\frac{(1+\gamma(\psi)+\alpha\psi)^{1.5069}}{\psi^{0.266053\nu-0.56165}}\,.$ (2.6)
There is a minimum value of $\psi$ for which Eq. (2.6) has solutions. Below this, the object cannot burn hydrogen stably because the core temperatures and densities are too low to sustain a sufficient nuclear burning rate to balance the surface luminosity. This boundary corresponds to the minimum mass for hydrogen burning (MMHB). Stars heavier than this are red dwarfs while lighter objects are brown dwarfs. In our model, the specific value is $M_{\rm MMHB}=0.075{\rm M}_{\odot}$, consistent with the literature e.g., [91, 88].
@@
Under the assumptions above, the DM heat injection rate is given at any point of the stellar evolution by [81, 100, 34]
@@
$\begin{split}L_{\rm DM}&=m_{\rm DM}\pi R^{2}\Phi,\quad\text{with}\\ \Phi&=\sqrt{\frac{8}{3\pi}}\frac{\rho_{\rm DM}v_{\rm DM}f_{\rm cap}}{m_{\rm DM}}\left(1+\frac{3}{2}\frac{v_{\rm esc}^{2}}{v_{\rm DM}^{2}}\right)\end{split}$ (2.7)
@@
where $v_{\rm esc}^{2}=2GM/R$ is the escape velocity, $\rho_{\rm DM}$ is the ambient DM density, $v_{\rm DM}$ is circular velocity of galactic DM, $f_{\rm cap}$ is the fraction of DM that is captured, and $m_{\rm DM}$ is the DM mass; the term in the brackets accounts for gravitational focusing. Throughout this work, we fix $v_{\rm DM}=50$ km/s as a conservative benchmark for the GC [101, 102]. These equations have an important consequence: the DM luminosity is independent of the DM mass. Since the properties of the object are determined solely by energy conservation, the characteristics of objects supported partially or fully by DM annihilation are similarly DM mass-independent. Similarly, our results are insensitive to the details of the DM profile because the total luminosity, found by integrating over this, is determined by the star’s radius and the DM flux $\Phi$. This implies that scenarios where there is a substantial amount of DM at the star’s surface [103] are also covered by our analysis.
We will set the fraction of dark matter captured, $f_{\rm cap}$, to unity in what follows. This translates into an assumption about mass and cross section which can be read of from Fig. S5 in [81]. However, we note that the observational signals which we predict can be observable when only a subfraction of dark matter is captured i.e., $f_{\rm cap}<1$ because the dark matter density near the galactic center exceeds the local value by several orders of magnitude. In our plots below, we show results for varying $\rho_{\rm DM}$ assuming $f_{\rm cap}=1$, however, they simply rescale as $\rho_{\rm DM}f_{\rm cap}$ for scenarios where dark matter capture is not 100% efficient. Another scenario that is accessible in high dark matter density environments is a setup where a subfraction of dark matter exhibits large elastic cross sections as expected in models with composite dark matter candidates [104, 105]. Here, our results apply by replacing the capture fraction $f_{\rm cap}$ of dark matter in the object by the total fraction of the strongly interacting dark matter sub-component in the halo $f_{\rm strong}$. This fraction can be sizable depending on the dark matter mass, and is largely untested by terrestrial searches, see for example Ref. [106].
@@
To understand the differences between hydrogen and DM burning, it is enlightening to consider a star supported entirely by DM annihilation, i.e., $L_{\rm DM}=L_{\rm surf}$. Using equation (A.5) for the radius in (2.7), we can find an analytic solution when gravitational focusing is neglected:
@@
$\begin{split}M_{0}=&8.58\times 10^{-7}\times\frac{(1+\gamma(\psi)+\alpha\psi)^{1.19895}}{\psi^{1.4993\nu}}\\ &\times\left(\frac{\rho_{\rm DM}}{\rm GeV\,cm^{-3}}\frac{v_{\rm DM}}{50\rm km\,s^{-1}}\frac{f_{\rm cap}}{b_{1}^{2.857}}\right)^{0.52475}.\end{split}$ (2.8)
@@
In contrast to the equivalent formula for hydrogen-supported objects (2.6), Eq. (2.8) has no minimum — DM can support structures of arbitrarily small masses. This is because, unlike hydrogen burning, the DM burning rate does not depend on the stellar properties — DM burning is ever-present. Including gravitational focusing does not alter this conclusion and in fact produces stronger effects. It is included in our numerical studies below.
@@
3 Time Evolution and Stability
To study the time evolution that leads to the formation of dark dwarfs, we use the cooling model introduced in [107, 91]. Using the first law of thermodynamics, the equation for the energy of a contracting star is
$T\frac{\mathrm{d}S}{\mathrm{d}t}=\varepsilon-\frac{\partial L_{\rm surf}}{\partial M}$ (3.1)
where $S$ is the entropy per unit mass and $\varepsilon$ is the energy generation rate per unit mass from burning. Integrating this equation, employing the equation of state for molecular hydrogen, and using the EOS and properties of $n=3/2$ polytropes one finds [88]
$\frac{\mathrm{d}\psi}{\mathrm{d}t}=\frac{\bar{\mu}}{1.5\mu_{e}^{\frac{8}{3}}E_{0}}M_{0}^{{-7/3}}(1+\gamma(\psi)+\alpha\psi)^{2}(L_{\rm burn}-L_{\rm surf}),$ (3.2)
@@ 3 Time Evolution and Stability
$\frac{1}{\bar{\mu}}=\frac{1}{{\mu}}+\frac{3X_{H^{+}}(1-X_{H^{+}})}{2(2-X_{H^{+}})}$ (3.3)
@@
with $X_{H^{+}}$ the mass fraction of ionized hydrogen; $\bar{\mu}=1.02$ for parameter point D, see the Supplementary Material. In the absence of any burning, equation (3.2) implies that $\psi$ will decrease from its initial value i.e., that over time $k_{B}T$ will decrease relative to the Fermi energy $E_{F}$ (see equation (A.3)) and the star will be increasingly supported by degeneracy pressure. In the presence of burning, there is a steady-state when the right hand side of equation (3.2) is zero i.e., when $L_{\rm burn}=L_{\rm surf}$. These are the solutions we derived above. We thus expect that any initial configuration will evolve towards this steady-state, reaching it at sufficiently late times.
@@
The stability of steady-state solutions can be determined as follows. Letting $\psi_{0}$ be the value of $\psi$ where the steady-state is achieved, we can write $\psi(t)=\psi_{0}+\delta\psi(t)$ and Taylor-expand equation (3.2) to find $\dot{\delta\psi}=f(\psi_{0})\delta\psi+\mathcal{O}(\delta\psi^{2})$, where $f(\psi_{0})$ is the derivative of the right hand side of (3.2) with respect to $\psi$ and evaluated at $\psi_{0}$. If $f(\psi_{0})<0$ then the equilibrium is stable. The resulting expressions are long and not informative so we do not give them here. They can be found in our accompanying  code. Using the same code, we found that dark dwarfs are always stable. This is true for objects supported by a combination of H-burning and DM burning, or objects supported solely by DM burning. In contrast, only the branch of SM stars (solely H-burning) with $\psi>\psi_{\rm min}$ is stable.
@@
Figure 3: Time evolution of important stellar properties. The continuous lines correspond to objects evolving in the presence of DM annihilation while dashed lines indicate the SM predictions. Four representative stellar masses indicated in the top panel are shown. In the SM, the lighter two objects evolve to become brown dwarfs, and the heavier objects evolve to become red dwarfs, as can be observed by the continuous decrease versus stabilizing of the luminosity at late times.
@@
Figure 4: Time evolution of the core temperatures of three objects with different masses with (solid lines) and without (dashed lines) DM energy injection. Coulomb corrections to the pressure are included. In the SM, the lighter objects evolve to become a brown dwarf, and the heavier objects evolve to become red dwarfs.
@@
In Fig. 3 we show properties of objects in the presence of DM annihilation as a function of time found by solving Eq. (3.2). At early times, the evolution of the objects are dominated by gravitational collapse so have similar properties. At later times, several differences between SM objects and those with some amount of DM burning are evident. SM brown dwarfs lighter than the MMHB cool continuously, contracting and becoming dimmer. SM objects heavier than the MMHB cool and evolve to red dwarfs where they are supported by H-burning. In contrast, in the presence of DM burning lighter objects cool similarly to brown dwarfs until they reach the steady-state dark dwarf solution with constant core temperature, radius, and brightness. The time to reach the dark dwarf state is shorter for larger ambient DM densities. The MMHB is also DM density dependent.
@@
4 Lithium burning
The lithium test is a primary method for confirming that an object is a brown dwarf [108]. Observing lithium lines in stellar spectra allows astronomers to trace the core temperature history of young stars and brown dwarfs, distinguishing between different evolutionary epochs. Therefore, deviations from the SM expectation of the lithium-7 abundance may provide a method for detecting dark dwarfs.